OpenStudy (anonymous):

sin(2cos^-1 4/5)= step by step

OpenStudy (agl202):

Do you know that cos^-1(4/5) = sin^-1(3/5) otherwise draw a diagram and convenience yourself!!! Next use Sin2A = 2sinAcosA so the given expression is 2sincos^-1 (4/5)coscos^-1 (4/5) = 2sinsin^-1 (3/5)coscos^-1 (4/5) =2(3/5)(4/5) =24/25

OpenStudy (anonymous):

awesome. thank you!!!